# Localized Deviations from the CO–PAH Relation in PHANGS–JWST Galaxies: Faint PAH Emission or Elevated CO Emissivity?

### v1

@@ Appendix A Full list of galaxies hosting CO/PAH outliers
Table 1 lists physical properties of 20 galaxies hosting outlier regions, together with the fraction of outlier pixels relative to the entire CO-emitting area. From Figures 6 to 11, we show all the galaxies studied in R. Chown et al. (2025) that host regions with abnormally high CO-to-PAH ratios. When available, three-color composite HST images from the PHANGS–HST survey (J. C. Lee et al., 2022) are also included for some galaxies. Red contours highlight abnormal regions, whereas their null test regions are identified in black.
@@ Appendix B CO velocity dispersions in individual galaxies
Figure 12 compares CO velocity dispersion () as a function of molecular gas surface density () between abnormal and null-test regions for all 20 galaxies hosting outlier regions.

### v2

@@ Appendix A Full list of galaxies hosting CO/PAH outliers
Table 1 lists physical properties of 20 galaxies hosting outlier regions, together with the fraction of outlier pixels relative to the entire CO-emitting area. From Figures 6 to 9, we show all the galaxies studied in R. Chown et al. (2025a) that host regions with abnormally high CO-to-PAH ratios. When available, three-color composite HST images from the PHANGS–HST survey (J. C. Lee et al., 2022) are also included for some galaxies. Red contours highlight outlier regions, whereas their null-test regions are identified in black.
@@ Appendix B CO velocity dispersions in individual galaxies
Figure 12 compares CO velocity dispersion () as a function of molecular gas surface density () between outlier and null-test regions for all 20 galaxies hosting outlier regions.